OpenStudy (anonymous):

the function f(x) varies directly with x and f(x)=3 when x=12 what is the value of f(x) when x=40

OpenStudy (ribhu):

f(x) = kx 3 = k * 12 k = 1/4 so f(x) = x/4 so when x = 40, f(x) = 10